### v1

# Towards Multi-modal Graph Large Language Model

Abstract
Multi-modal graphs, which integrate diverse multi-modal features and relations, are ubiquitous in real-world applications. However, existing multi-modal graph learning methods are typically trained from scratch for specific graph data and tasks, failing to generalize across various multi-modal graph data and tasks. To bridge this gap, we explore the potential of Multi-modal Graph Large Language Models (MG-LLM) to unify and generalize across diverse multi-modal graph data and tasks. We propose a unified framework of multi-modal graph data, task, and model, discovering the inherent multi-granularity and multi-scale characteristics in multi-modal graphs. Specifically, we present five key desired characteristics for MG-LLM: 1) unified space for multi-modal structures and attributes, 2) capability of handling diverse multi-modal graph tasks, 3) multi-modal graph in-context learning, 4) multi-modal graph interaction with natural language, and 5) multi-modal graph reasoning. We then elaborate on the key challenges, review related works, and highlight promising future research directions towards realizing these ambitious characteristics. Finally, we summarize existing multi-modal graph datasets pertinent for model training. We believe this paper can contribute to the ongoing advancement of the research towards MG-LLM for generalization across multi-modal graph data and tasks.
Index Terms:
Multi-modal Graph, Large Language Model, Foundation Model, Graph Machine Learning, Multi-modality
I Introduction
Multi-modal graphs, which integrate features from diverse modalities such as text, image, audio and video, as well as capture the complex intra-model and inter-modal relations, are becoming increasingly ubiquitous in real-world applications. From social networks [bhattacharyya2024heterogeneous] and e-commerce [wang2023fashionklip] platforms to scientific discovery in biomedicine and materials science [marini2024multimodal, ye2024construction], these complex data structures offer a richer, more holistic representation of interconnected entities than traditional unimodal graphs, which unlock more opportunities for advanced analytics, reasoning, and generation capabilities over multimodal information.
Figure 1: Key characteristics towards Multi-modal Graph Large Language Models (MG-LLM).
However, multi-modal graph learning currently faces a significant issue: existing methods are predominantly designed for specific tasks on particular types of graphs. This specialization often limits their applicability, preventing them from generalizing effectively across the vast diversity of multi-modal graph data and tasks encountered in practice. This lack of universality necessitates constant redesign and retraining for new scenarios, hindering the development of truly versatile and scalable solutions.
To bridge this gap, we explore the potential of Multi-modal Graph Large Language Models (MG-LLM). Inspired by the remarkable success of large language models (LLMs) in unifying diverse natural language tasks, we propose that MG-LLM can serve as a powerful paradigm to unify and generalize across the complex landscape of multi-modal graph data and tasks. Our exploration begins by establishing a unified framework for multi-modal graph data, tasks, and models, which uncovers the inherent characteristics of multi-modal graphs, i.e., multi-granularity and multi-scale.
Specifically, we highlight that multi-modal graphs inherently exhibit multi-granularity, organizing information from fine-grained features like pixels and words to coarse-grained concepts such as entire images or documents, along with diverse structural complexities. This leads to multi-scale characteristics in multi-modal graph tasks, where inputs and outputs can vary dramatically in their scope, from individual nodes to entire graph structures.
Building upon this foundational understanding, we articulate five key desired characteristics towards MG-LLM:
  • Unified Space for Multi-modal Structures and Attributes: the ability to align and represent diverse multi-modal features and relations within a single unified embedding space, capable of handling highly irregular and continuous information.
  • Ability of Handling Diverse Multi-modal Graph Tasks: the capacity to frame and solve all multi-modal graph tasks, from traditional discriminative problems like node classification to emerging generative tasks such as multi-modal content generation, under a unified generative modeling paradigm.
  • Multi-modal Graph In-context Learning: the capability of performing novel tasks by leveraging a limited number of multi-modal graph examples provided directly within the prompt, without requiring explicit model fine-tuning.
  • Multi-modal Graph Interaction with Natural Language: the possibility of enabling users to query, edit, generate, and reason about complex multi-modal graph-structured knowledge using intuitive natural language, bridging the gap between human language and structured data.
  • Multi-modal Graph Reasoning: the proficiency in performing complex multi-hop, cross-modal reasoning, including analogical inference, by seamlessly combining information from various modalities and relational structures.
While the vision of MG-LLMis ambitious, realizing these characteristics presents significant challenges, ranging from developing unified multi-modal graph vocabularies and tokenization schemes to multi-modal graph architectures capable of large-scale pretraining. This paper delves into these key challenges, reviews existing research that moves towards this paradigm, and outlines promising future research directions to accelerate the development of MG-LLM for generalizing across diverse multi-modal graph data and tasks. Finally, we summarize existing multi-modal graph datasets that could be useful for the training and evaluation of such models. This work aims to foster progress towards a new era of multi-modal graph intelligence.
Our main contributions are summarized as follows:
  • We explore the potential of Multi-modal Graph Large Language Models (MG-LLM) to unify and generalize across diverse multi-modal graph data and tasks, aiming for universal generalization across diverse multi-modal graph data and tasks. We systematically discuss the potential of MG-LLM, for the first time, to the best of our knowledge.
  • We present a unified framework for understanding multi-modal graph data, tasks, and models, highlighting their inherent multi-granularity and multi-scale characteristics for designing MG-LLM.
  • We propose five essential characteristics that MG-LLM should possess, coupled with detailed discussions of challenges and future directions, thereby setting a clear research roadmap. We also summarize existing multi-modal graph datasets and tasks for developing MG-LLM.
The domain of multi-modal learning on graphs is rapidly advancing. We summarize the differences of this work with two most related papers as follows: 1) Ektefaie et al. [ektefaie2023multimodal] focus on the challenges and methodologies of combining modalities within graph-based machine learning, while we take a distinct perspective by envisioning MG-LLM as a unifying paradigm, leveraging the power of large language models to address the limitations of task-specific MGNNs. 2) Peng et al. [peng2024learningmultimodalgraphssurvey] provide a comprehensive survey of existing techniques, data types, and applications within multi-modal graph learning, while we focus on introducing a novel conceptual framework and put forth the vision of MG-LLM. By identifying the inherent characteristics of multi-modal graphs and the desired characteristics for MG-LLM, we chart a new research roadmap towards developing multi-modal graph large language models, thereby complementing surveys of the existing state-of-the-art by looking towards a next-generation paradigm.
II Towards a Unified View of Multi-modal Graph Data, Task, and Model
In this section, we introduce a unified framework of multi-modal graph data, task, and model, and remarks on the inherent characteristics in multi-modal graphs, serving as a foundation to discuss the desired characteristics of MG-LLM. The overall framework is shown in Figure 2.
II-A Unified Formulation of Multi-modal Graph Data
In this section, we define multi-modal graphs, extending standard graphs with diverse node and edge modalities. Then we outline three decomposable types (feature-, node-, graph-level), and their versatility in representing various data forms, from single instances to full datasets. Moreover, we give remarks on the challenges of indecomposability and multi-granularity for building effective MG-LLM.
Graph
A graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ consists of a finite set of vertices $\mathcal{V}=\{v_{1},v_{2},\dots,v_{n}\}$ and a set of edges $\mathcal{E}\subseteq\mathcal{V}\times\mathcal{V}$, each edge being an ordered pair of vertices denoting a directed relation between them.
Figure 2: Unified view of multimodal graph data, tasks and model towards MG-LLM.
Multi-modal Graph
A modality is a distinct type or source of information associated with nodes or edges. Let =$\mathcal{M}=\{1,2,\dots,M\}$ denote the set of all modalities. For each $m\in\mathcal{M}$, we define a mapping $F_{m}$ from the modality-specific node feature space $\mathcal{V}$ to a shared representation space $\mathcal{X}$, such that $F_{m}:\mathcal{V}\to\mathcal{X}$. We define a mapping set =$\mathcal{F}=\{F_{m}\}_{m=1}^{M}$ for all $M$ modalities. The space $\mathcal{X}$ serves as a unified embedding space for all modalities. Similarly, we define a mapping ${F}_{m}$ from the modality-specific edge feature space $\mathcal{E}$ to the shared representation space $\mathcal{X}$, such that $\mathcal{F}_{m}:\mathcal{E}\to\mathcal{X}$. For simplicity, we reuse ${F}_{m}$ as the modality-$m$ map for both nodes and edges. We likewise leave out multimodal edges in the formulation, even though extending them would be straightforward. The features in different modalities could be texts, images, audios, videos, etc. A multi-modal graph can be defined by the quadruple $\mathcal{G}=(\mathcal{V},\mathcal{E},\mathcal{F},\mathcal{M})$.
Special Cases of Decomposable Multi-modal Graphs
By instantiation of the modality set and feature mapping, we could obtain several classic types of multi-modal graphs which are ubiquitous in real-world applications [peng2024learningmultimodalgraphssurvey]. These types of multi-modal graphs share the same assumption that they can be decomposed by modality from different perspectives, i.e., feature, node, and graph:
  • Feature-level Multi-modal Graph, where the features of nodes or edges come from different modalities, i.e., $\mathcal{G}=\bigcup_{m=1}^{M}\mathcal{G}_{m}=\bigcup_{m=1}^{M}(\mathcal{V},\mathcal{E},\mathcal{F}_{m},\{m\}),$ where $m$ representing one modality in $M$ modalities. The feature of node $v$ could be represented as $\mathbf{x}(v)=\bigoplus_{m=1}^{M}\mathcal{F}_{m}(v)$. For example, on an e-commerce product graph [zhu2025mosaicmodalitiescomprehensivebenchmark], each node has the feature of product title and image, i.e., $\mathbf{x}(v)=\bigoplus(\mathcal{F}_{{\text{text}}}(v),\mathcal{F}_{{\text{image}}}(v))$.
  • Node-level Multi-modal Graph, where the nodes or edges come from different modalities, while each node or edge has unimodal features, i.e., $\mathcal{G}=\bigcup_{m=1}^{M}\mathcal{G}_{m}=\bigcup_{m=1}^{M}(\mathcal{V}_{m},\mathcal{E},\mathcal{F}_{m},\{m\}),$ where $m$ representing one modality in $M$ modalities, and $\mathcal{V}_{i}\bigcap\mathcal{V}_{j}=\emptyset$. For example, on a multimodal knowledge base [10.1007/978-3-030-41407-8_9], each node might be either an image or a textual description.
  • Graph-level Multi-modal Graph, where the graphs come from different modalities, while each graph has unimodal features, i.e., $\mathcal{G}=\bigcup_{m=1}^{M}\mathcal{G}_{m}=\bigcup_{m=1}^{M}(\mathcal{V}_{m},\mathcal{E}_{m},\mathcal{X}_{m},\{m\}),$ where $m$ representing one modality in $M$ modalities, and $\mathcal{E}_{i}\bigcap\mathcal{E}_{j}=\emptyset$. For example, on a multi-modal question answering graph [8953451], we may have a graph with images, a graph with texts, etc.
Remark 1.
(Indecomposable Characteristics) Although practitioners may model their data with the aforementioned decomposable multi-modal graphs for convenience, most multi-modal graphs in real-world scenarios, with nodes and edges having features from various modalities, may not be easily decomposable to several uni-modal subgraphs. For instance, in a multi-modal graph where the text node says ‘The Transformer was incredible!’, the image node shows Optimus Prime (a central robot character from the Transformers movie series), and the knowledge node links to the movie Transformers, only joint reasoning over all three nodes can resolve the ambiguity and correctly interpret ‘Transformer’ as a film character rather than a neural network architecture. Due to the indecomposable characteristics, established multi-modal fusion techniques in other multi-modal fields [zhao2024deep] may fail to flexibly solve multi-modal graph problems, calling for the need of native modeling of multi-modal graph data in multi-modal graph large language models.
Special Cases of Multi-modal Graph Instances
The versatility of multi-modal graphs allows them to represent not only complex inter-modal relationships but also instances or datasets composed of single or multiple modalities as special cases, e.g., instances of texts, images, audios, videos, etc, or pairs of image-captions, text-audios, etc. Here are examples of representing single-modal instances:
  • a text sequence can be represented by a multi-modal graph with a single text-attributed node, i.e., $\mathcal{G}=\left(\mathcal{V}=\{v\},\mathcal{E}=\emptyset,\{F_{\text{text}}\},\{\text{text}\}\right)$, where $F_{\text{text}}$ is the text feature mapping function.
  • a text sequence, more granularly, can be represented by a multi-modal graph where each word is a node and sequential or semantic connections form edges, i.e., $\mathcal{G}=\left(\mathcal{V},\mathcal{E},\{F_{\text{word}}\},\{\text{word}\}\right)$, where $\mathcal{V}=\{v_{1},\dots,v_{L}\},\mathcal{E}=\{(v_{i},v_{i+1})\}_{i=1}^{L-1}$, $L$ is the length of the text sequence, and $F_{\text{word}}$ is the word feature mapping function.
  • an image can be represented as a multi-modal graph where pixels are nodes and their grid-like inter-connections form edges, i.e., an $H\times W$ image can be $\mathcal{G}=\left(\mathcal{V}=\{v_{ij}\}_{i=1,j=1}^{H,W},\mathcal{E}_{\text{grid}},\{F_{\text{pixel}}\},\{\text{pixel}\}\right)$, where $\mathcal{E}_{\text{grid}}$ represents grid-like inter-connections, and $F_{\text{pixel}}$ is the pixel feature mapping function.
Beyond individual instances, multi-modal graphs can efficiently represent entire datasets. Here are some examples.
  • A batch of images can be represented as a multi-modal graph with several image-attributed nodes without any edges, i.e., $\mathcal{G}=\left(\mathcal{V}=\{v_{1},\dots,v_{K}\},\mathcal{E}=\emptyset,\{F_{\text{image}}\},\{\text{image}\}\right)$, where $F_{\text{image}}$ is the image feature mapping function, and $K$ is the number of images.
[…]
This ability to abstract various data forms into a unified graph structure underscores the expressive power of multi-modal graphs.